# MEMIT-Merge: Addressing MEMIT’s Key-Value Conflicts in Same-Subject Batch Editing for LLMs

### v2

@@ 5.4 Results with Varying Batch Sizes
As can be seen in Fig. 4, when the subjects of the editing knowledge in the edit batch are the same, the standard edit success rate plummets at a batch size of 2, whereas MEMIT-Merge is able to maintain a much higher success rate, with a significantly smaller decline compared to MEMIT. This also confirms the effectiveness of our method.
In the case of distinct subjects, the editing capability of both MEMIT and MEMIT-Merge does not exhibit a significant decline even at a batch size of 100, which is consistent with our previous analysis. The results of other editing methods are given in App. C.
@@ 6 Conclusion
This paper identifies the issue of significant performance degradation in MEMIT when a batch contains knowledge sharing the same subject during batch editing. We analyze the reason based on the MEMIT framework, and develop the metric $AKD$ to further analyze the cause of this phenomenon. This study identifies a critical flaw in MEMIT’s batch editing: parameter update conflicts arising from identical keys requiring divergent values in same-subject scenarios. Our proposed MEMIT-Merge resolves this through key-wise value merging, significantly improving same-subject editing success while maintaining original performance on distinct-subject cases. These findings advance mass-editing techniques for evolving LLM knowledge bases.
@@ Appendix A Details of Constructing Same-Subject and Distinct-Subject Data
Subsequently, following the dataset metrics in Meng et al. (2022), we add two types of questions: specificity and paraphrase. For paraphrase questions, we use the same knowledge triples as the edit sentences, but with a different template format: “The name of the relation of subject is object.”. For specificity, there are two types of questions. One is completely unrelated knowledge, for which we use the prompt “The capital city of America is”. The other type has the same relation as the edited knowledge but a different subject. For example, if the edited knowledge is (John, father, Bob), a specificity question could be (Paul, father, Eugène).
Appendix B Diverse $AKD$ Dataset
@@
Figure 7: Editing same-subject dataset using Llama-3-8B-Instruct with four editing methods.
Appendix C More Results with Varying Batch Sizes
Here in Fig. 6 and Fig. 7 we demonstrate some more results about editing same subject batch with varying batch sizes.
It shows clearly that MEMIT-based methods suffers from same subject issue, while methods like FT do not.

### v3

@@ 4.4 Results with Varying Batch Sizes
As can be seen in Fig. 3, when the subjects of the editing knowledge in the edit batch are the same, the standard edit success rate plummets at a batch size of 2, whereas MEMIT-Merge is able to maintain a much higher success rate, with a significantly smaller decline compared to MEMIT. This also confirms the effectiveness of our method. The results of other models are given in App. E
In the case of distinct subjects, the editing capability of both MEMIT and MEMIT-Merge does not exhibit a significant decline even at a batch size of 100, which is consistent with our previous analysis.
It can also be observed that even when utilizing MEMIT-Merge, the outcomes for same-subject scenarios do not yet align with those achieved in distinct-subject scenarios. Although we have resolved the conflicts between key-value pairs, the capability of a value vector is not unlimited and may cause a new issue. We have enabled the value to match more relations to target objects, but performance may decline with very large batches, as a single value vector cannot handle an infinite number of associations. This is why MEMIT-Merge performs well with small batches but worsens as the batch size increases. Nonetheless, real-world editing typically involves a small number of same-subject edits (usually fewer than 10), and each subject has a limited set of properties. Given this, our method effectively addresses same-subject cases without notable performance loss.
The results of other editing methods are given in App. F.
@@ 5 Conclusion
This paper identifies the issue of significant performance degradation in MEMIT when a batch contains knowledge sharing the same subject during batch editing. This is fundamentally caused by parameter update conflicts arising from identical keys requiring divergent values in the same-subject scenarios. Our proposed MEMIT-Merge resolves this and significantly improves same-subject edit performance while maintaining original performance on distinct-subject cases. These findings advance mass-editing techniques for evolving LLM knowledge bases.
Limitations
While this study provides insight into the same-subject issues within the MEMIT-based method, several limitations should be recognized. First, all knowledge triples are restricted to person-related entities, leaving the generalization to other subject types (such as locations or organizations) untested. While our theoretical framework suggests that subject type should not fundamentally alter the conclusions, empirical validation across diverse categories remains necessary. Second, the experiments focus solely on lexical-level subject distinctions; potential effects of semantic similarity in embedding space were not explored. Future work could extend this investigation by incorporating larger datasets, multi-type knowledge triples, and embedding-space analyses to further validate the theoretical predictions.
Furthermore, although we have resolved the conflicts between key-value pairs, the capability of value vector is not unlimited and may cause a new issue. The core factor, ultimately, is that the key vector at the position of the subject’s last token does not contain information about the relation. Our approach is a solution that does not alter the position of the edited fact tokens. To thoroughly resolve this problem, it may be necessary to avoid using the subject’s last token and instead use the last token or other positions that can incorporate the’relation’ information for knowledge editing.
Acknowledgments
This work was supported by the Natural Science Foundation of China (No. 62476134).
@@ Appendix A Details of Constructing Same-Subject and Distinct-Subject Data
Subsequently, following the dataset metrics in Meng et al. (2022), we add two types of questions: specificity and paraphrase. For paraphrase questions, we use the same knowledge triples as the edit sentences, but with a different template format: “The name of the relation of subject is object.” For specificity, there are two types of questions. One is completely unrelated knowledge, for which we use the prompt “The capital city of America is.” The other type has the same relation as the edited knowledge but a different subject. For example, if the edited knowledge is (John, father, Bob), a specificity question could be (Paul, father, Eugène).
Appendix B Further Analysis of Cause
[…]
We compute the $AKD$ for all layers of the model at the subject’s last token position. As the degree of subject variation increases across sentences, the $AKD$ value proportionally rises. Conversely, when all sentences share identical subjects, the $AKD$ value remains constant at 0.
We construct sentence batches using predefined templates, where batches sharing the same template exhibited similar $AKD$ values, while distinct templates yielded significantly different $AKD$ measurements. The specific templates and corresponding $AKD$ values are detailed in App. C. For experimental validation, we select three $AKD$ groups (0, 10, 25) and conduct editing tests using Qwen2.5-1.5B-Instruct. As shown in Fig. 4, where $AKD$ values are computed using keys from MEMIT’s final editing layer, the results demonstrate an inverse relationship: lower $AKD$ values correspond to reduced editing success rates. This pattern remains consistent across other $AKD$ values, establishing a statistically significant negative correlation between $AKD$ and editing efficacy.
Figure 4: Datasets with different $AKD$ values and the results of edit efficacy. The lower the AKDAKD value, the more severe the decline in edit capability.
Appendix C Diverse $AKD$ Dataset
@@
Appendix D Results of Other Models At Batch Size 10
Tab. 3 shows results using Llama-3-8B-Instruct, GPT-J-6B and Qwen2.5-7B-Instruct at batch size 10. It confirms the same-subject issue on MEMIT and the advantage of our MEMIT-Merge across different LLM model sizes and architectures.
Model Dataset Method Efficacy Paraphrase Specificity
Llama-3-8B-Instruct same-dataset FT 0.68 0.63 0.60
MEMIT 0.51 0.40 0.99
MEMIT-Merge 1.00 0.78 0.99
distinct-dataset FT 0.69 0.65 0.59
MEMIT 1.00 0.90 0.95
MEMIT-Merge 1.00 0.93 0.96
Qwen2.5-7B-Instruct same-dataset FT 0.31 0.23 0.98
MEMIT 0.43 0.37 1.00
MEMIT-Merge 0.99 0.49 0.99
distinct-dataset FT 0.26 0.23 0.98
MEMIT 1.00 0.82 0.98
MEMIT-Merge 1.00 0.86 0.97
GPT-J-6B same-dataset FT 0.73 0.57 0.32
MEMIT 0.34 0.25 1.00
MEMIT-Merge 1.00 0.35 1.00
distinct-dataset FT 0.74 0.69 0.35
MEMIT 1.00 0.77 0.99
MEMIT-Merge 1.00 0.77 0.99
Table 3: results of models in various size and architecture when batch size is 10.
Appendix E Results with Varying Batch Sizes of Other Models
Appendix F Results with Varying Batch Sizes of other methods
Here in Fig. 5 we demonstrate some more results about editing same subject batch with varying batch sizes.
It shows clearly that MEMIT-based methods suffer from the same subject issue, while methods like FT do not.
Appendix G Related Work
Knowledge editing techniques for large language models (LLMs) primarily fall into two paradigms: non-parametric approaches that preserve original parameters and parametric methods that directly modify model weights. Parametric approaches, while effective for targeted updates, often introduce uncontrolled parameter perturbations that adversely affect unrelated knowledge, a challenge addressed through various constraint mechanisms. The parametric category features two dominant subclasses: one is “meta-learning based methods”, such as MEND Mitchell et al. (2022) and MALMEN Tan et al. (2024) that train meta-networks using carefully designed datasets containing both unrelated knowledge samples and paraphrased sentences, aiming to enhance generalization while minimizing collateral damage. Another is locate-and-edit Methods, which includes techniques such as Knowledge Neuron Dai et al. (2022), identifying critical knowledge storage locations before executing precise edits. ROME Meng et al. (2022) extended this by incorporating knowledge preservation terms in its optimization objective to maintain model integrity.
Our work builds upon MEMIT Meng et al. (2023), a state-of-the-art locate-and-edit approach that enables batch knowledge editing through MLP layer modifications. Building on MEMIT, many recent methods have made modifications to parameter update methods during editing or to the architecture and location of the edits. PMET Li et al. (2024) incorporated the output of the attention layer in the calculation of parameter updates. AlphaEdit Fang et al. (2025) improved upon MEMIT’s parameter matrix update method by projecting the update matrix into the null space of the original knowledge to mitigate interference with unrelated knowledge. UNKE Deng et al. (2024) extended structured knowledge editing to unstructured editing.